. During the Great Recession, U.S investors suffered a huge loss in business confidence once the stock market crashed in 2008. Consider the model below: Y = 1650; C = 200 + 0.6(Y – T); G = 250; I = 400 – 2000r; T= 150 i. Suppose that taxes are cut by 10%, so that the new level of taxes is T= 135. Find the new equilibrium r, I, and C. How much investment is crowded out by the tax cut? iii. Compute the new public savings, private savings, and national savings after the tax cut.
Answer #1

Y = cth tI 1650 = 200 +0.6 (1650-135) +250 troon 2000r. 109= 20000 8= 0.0545 I= 400 – 2000 (0.0545) = 291 (= 200 to 6 (165o -